OpenStudy (anonymous):

The lengths of the legs of a right triangle are 8 inches and twelve inches. What is the exact length of the hypotenuse?

OpenStudy (johnweldon1993):

Remember that for a right triangle you can use Pythagoreas theorem! It states that the sum of both legs squared = the hypotenuse squared....written in an equation looks like \[a^2 + b^2 = c^2\] You have the 2 legs....8 and 12.....and you want to solve for the hypotenuse....so you plug in these numbers \[8^2 + 12^2 = c^2\] or to find the exact length you do \[c= \sqrt{8^2 + 12^2}\] what do you get?